OpenStudy (anonymous):

Tank is shape of right circular cylinder with radius of 3m and height of 10m axis is horizontal. tank is full of gasoline with density of 680 kg/m^3 calculate work to empty tank through a hole at the top of one of the circular ends.

OpenStudy (perl):

you can integrate work here

OpenStudy (anonymous):

yeah is my F= 680(9pi)?

OpenStudy (anonymous):

is r constant at 3m?

OpenStudy (perl):

we can cut the cylinder into slices

OpenStudy (anonymous):

because when i was doing the sphere the r changes and i had to use x=sqrt(stuff)

OpenStudy (perl):

the i'th slice

OpenStudy (perl):

what is the work to move the ith slice up a distance of 10 - y

OpenStudy (anonymous):

yeah i get that, but i don't know the Force

OpenStudy (perl):

force is weight

OpenStudy (anonymous):

usually force is equal to the density times the volume correct?

OpenStudy (perl):

density = mass / volume. volume * density = mass

OpenStudy (perl):

then mass * g (gravity constant) = weight

OpenStudy (perl):

the weight of the ith slice is mass * gravitational acceleration constant (g) but mass = volume * density weight = Volume * density * g pi * r^2 * dy * density * g

OpenStudy (anonymous):

what is the function of "r" in this problem?

OpenStudy (perl):

r = 3

OpenStudy (perl):

we use r in the formula for the slice of volume (it looks like a pancake)

OpenStudy (perl):

the work to empty a cylindrical tank of a liquid of density rho is the integral Integral {y = 0 , y = h} pi * r^2 * dy * rho * g * (h - y )

OpenStudy (perl):

here h = 10, rho = 680, r = 3 , g = 9.8 you might be able to factor out all the constants first

OpenStudy (anonymous):

this is for calculus not physics, we don't use g

OpenStudy (perl):

yes you have to use g, since you are given kg/m^3 for density

OpenStudy (anonymous):

volume should be: integral from 0 to 10 of 680(9)(pi)(10-y) dy?

OpenStudy (perl):

use a better estimate for g

OpenStudy (anonymous):

my 9 is from r^2

OpenStudy (perl):

oh you left out g then

OpenStudy (perl):

we can try it your way

OpenStudy (perl):

yes , integrate that

OpenStudy (perl):

you should get 50 for that

OpenStudy (perl):

integral 680(9)(pi)(10-y) = 680 * 9 * pi * 50 (but i still think this is incorrect)

OpenStudy (anonymous):

ok, i don't know about multiplying by g i thought that since density was given we didn't need to do that

OpenStudy (perl):

usually these calculus problems are given in units of pounds and feet , so you don't need to multiply by g . this problem is given in metric system, its different

OpenStudy (perl):

we are not given the weight density, thats the problem. we are given the 'mass' density.

OpenStudy (anonymous):

i see thank you

OpenStudy (perl):